# 1.27in1.27in1in1in

### v1

@@ 6.3. Type II Flops on Picard
Note that there is a small ambiguity here: there is no canonical correspondence between the $d_{i}$ and the $d_{i}^{\prime}$. Contrary to what the notation suggests, they are not canonically ordered, so that once we have chosen an ordering (which we will tacitly do), sending $d_{1}\mapsto d_{2}^{\prime}$ and $d_{2}\mapsto d_{1}^{\prime}$ would be an equally reasonable choice. We do not pay attention to this, since this issue only plays a role on the normalization and we are ultimately interested in the glued surfaces. Suppose for the sake of the argument that $Y^{\prime}$ is of class $T$ (and thus $Y$ is of class $P$). Then the curves $d_{i}^{\prime}$ for $i=1,2$ are glued to one another under the normalization map. While the combinatorial pushforward map on $\operatorname{Pic}(Y^{\prime\nu})$ to be defined below will depend on the numbering of the $d_{i}^{\prime}$, its restriction to $\operatorname{Pic}(Y^{\prime})$ will not. This is reflected graphically in the fact that the curves $d_{i}^{\prime}$ are exchanged under the graph automorphism of the augmented curve structure that fixes the upper row and the white $(-2)$-vertex in the lower row of Figure 5 and swaps the two leftmost vertices with the two rightmost vertices in the lower row.
@@
(6.7) $\phi_{c}(v):=\begin{cases}-v^{\prime}&\textrm{ if }v\in\{d_{1},d_{2}\}\\ v^{\prime}&\textrm{ otherwise.}\\ \end{cases}$
@@
Observe that $B_{1}+B_{2}\in\operatorname{Pic}(Y)$ and this is a square $(-2)$ class. As in Example 6.7, there is a unique divisor $D$ on ${\mathcal{Y}}$ that restricts to $D_{0}=B_{1}+B_{2}$ on the central fiber. We only have to show that $\phi_{c}(D_{0})=(\Phi_{*}D)_{0}$. We use Lemma 6.6 and the notation therein and write $D^{\prime}:=\Phi_{*}D$ and
@@
00 0 $=D.p_{*}F^{\prime}={\widetilde{D}}.F^{\prime}+\lambda E.F^{\prime}$
@@
00 0 $=D^{\prime}.q_{*}F={\widetilde{D}}.F+\lambda^{\prime}E.F$

### v2

@@ 6.3. Type II Flops on Picard
Note that there is a small ambiguity here: there is no canonical correspondence between the $d_{i}$ and the $d_{i}^{\prime}$. Contrary to what the notation suggests, they are not canonically ordered, so that once we have chosen an ordering (which we will tacitly do), sending $d_{1}\mapsto d_{2}^{\prime}$ and $d_{2}\mapsto d_{1}^{\prime}$ would be an equally reasonable choice. We do not pay attention to this, since this issue only plays a role on the normalization and we are ultimately interested in the glued surfaces. Suppose for the sake of the argument that $Y^{\prime}$ is of class $T$ (and thus $Y$ is of class $P$). Then the curves $d_{i}^{\prime}$ for $i=1,2$ are glued to one another under the normalization map. While the combinatorial pushforward map on $\operatorname{Pic}(Y^{\prime\nu})$ to be defined below will depend on the numbering of the $d_{i}^{\prime}$, its restriction to $\operatorname{Pic}(Y^{\prime})$ will not. This is reflected graphically in the fact that the curves $d_{i}^{\prime}$ are exchanged under the graph automorphism of the augmented curve structure that fixes the upper row and the white (2)(-2)-vertex in the lower row of Figure 5 and swaps the two leftmost vertices with the two rightmost vertices in the lower row.
@@
[…]
@@
Observe that $B_{1}+B_{2}\in\operatorname{Pic}(Y)$ and this is a square (2)(-2) class. As in Example 6.7, there is a unique divisor $D$ on $\mathcal{Y}$} that restricts to $D_{0}=B_{1}+B_{2}$ on the central fiber. We only have to show that $\phi_{c}(D_{0})=(\Phi_{*}D)_{0}$. We use Lemma 6.6 and the notation therein and write $D^{\prime}:=\Phi_{*}D$ and
@@
$0$ $=D.p_{*}F^{\prime}={\widetilde{D}}.F^{\prime}+\lambda E.F^{\prime}$
@@
$0$ $=D^{\prime}.q_{*}F={\widetilde{D}}.F+\lambda^{\prime}E.F$